OpenStudy (anonymous):

20. a bag contains 2 white marbles and 7 purble marbles. two marbles are drawn at random. one marble is drawn and is not replaced. then a second marble is drawn. 1. what is the probability of selecting a purple marble the a white marble? 2. What is the probability of selecting two white marbles? 3. Is there a greater chance of selecting two white marbles in a row or two purple marbles in a row? 21. Hallie is trying to win the grand prize on a game show. should she try her lucky by spinning a wheel with 6 equal sections labeled from 1 to 6

OpenStudy (anonymous):

and hope she gets a five or should she roll two two number cubes and she gets the same number on both cubes. explain

OpenStudy (anonymous):

20. 2 white, 7 purple total marbles = 9 probability of a white marble= 2/9 probability of a purple marble= 7/9

OpenStudy (anonymous):

thanks,what about 21?

OpenStudy (anonymous):

probability of getting a five on the spinner = 1/6 probability of getting 2 of the same numbers on dice = 1/12

OpenStudy (anonymous):

thanks dude

OpenStudy (anonymous):

no problem. :)